Species I has 2n = 8 chromosomes and species II ha
Answer #1

Species I: n1 = 4, Species II: n2=7:

I. Allotriploid: n1n1n1+n2n2+n2 = 4+4+4+7+7+7 = 12+21 = 33

II. Autotetraploid: 7x4 = 28

III. Trisomy: 2n+1 = 8+1 = 9

IV. Monosomy = 2n-1 = 14-1 = 13

V. Allotetraploid= 4x4 + 7x4 = 16+28 = 44